OpenStudy (anonymous):

8! over 3! x5! solve giving medals

OpenStudy (anonymous):

\[\frac{8!}{3! \times 5!}=\frac{8\times7\times6(5!)}{3\times2\times1 \times(5!)}\] \[=\frac{8\times7\times6}{3\times2\times1}=56\]

OpenStudy (anonymous):